OpenStudy (anonymous):

The length of a rectangle is 2 ft longer than its width. If the perimeter of the rectangle is 36ft, find its area

OpenStudy (anonymous):

l=w+2 36=2l+2w 36=2(w+2)+2w 36=2w+4+2w 36=4w+4 0=4w+4-36 -4w=4-36 divide everything by -4 w=-1+9 w=8 l=8+2=10 8*10=80 a=80ft^2